Bought service vehicle worth 200,000; paid 40% cash, while the balance in 4 equal monthly installments with 2% interest per month.
How much is the interest?

Answer by Theo(13342)

40% of 200,000 was paid in cash.
that left 60% to be financed.
200,000 * .6 = 120,000 that was financed.
this was to be paid in 4 equal monthly installments with 2% interest rate per month.

the financial calculator at C:\Users\dtheo\Pictures\G27.1041.53.jpg can be used to solve this.

there are formulas you can use as well, but the calculator is much easier and uses the same formulas.
if you need to use formulas, let me know and i'll send you the appropriate formula for this problem.

i used the calculator and came up with the following results.

the payment at the end of each month was 31,514.85.

there were 4 payments for a total payment of 126,059.4.

since the amount due was 120,000, then the difference was the interest on the loan.

that was equal to 6,059.4.

since the number of payments was small, you could actually calculate what happened month by month.

it works like this.
the amount of the loan was 120000
the payment at the end of each month was 31514.85 rounded to the nearest penny.

at the end of the first month, the amount owed became 120000 * 1.02 = 122400.
the payment was subtracted from that to get a remaining balance of 90885.15.

at the end of the second month, the amount owed became 90885.15 * 1.02 = 92702.853.
the payment was subtracted from that to get a remaining balance of 61188.003.

at the end of the third month, the amount owed became 61188.003 * 1.02 = 62411.76306.
the payment was subtracted from that to get a remaining balance of 30896.91306.

at the end of the fourth month, the amount owed became 30896.91306 * 1.02 = 31514.85132.
the payment was subtracted from that to get a remaining balance of .0013212.

without rounding of intermediate results, that amount would have been 0.
in fact, if you round the remaining balance to 2 decimal places, it becomes 0.

the total number of payments was 4.
31,514.85 * 4 = 126,059.40.
since the loan was for 120,000, then the interest paid was 6,059.40.

that's your solution.

here's what the calculator results look like.



inputs are everything except payment.
output is payment.